OpenStudy (iwanttogotostanford):

ACT MATH PREP Q: What is the slope-intercept for of 8x-y-6=0???

OpenStudy (iwanttogotostanford):

@Musicdude @

OpenStudy (anonymous):

so to get slope intercept you have to isolate y

OpenStudy (jdoe0001):

hmmm if you solve that for "y", what does it look like?

OpenStudy (anonymous):

so add 6 to the other side subtract 8x then change the signs because you divide by -1

OpenStudy (anonymous):

thats all

OpenStudy (iwanttogotostanford):

i got y=8x+6 I already know the answer, I just need to know how to do it not the answer

OpenStudy (jdoe0001):

hmmm check your signs again

OpenStudy (jdoe0001):

but that's quite close

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

y=8x-6

OpenStudy (anonymous):

the way you do this is by trying to get it to slope-intercept form which is y=mx+b

OpenStudy (jdoe0001):

\(8x-y-6=0\implies \begin{array}{llll} {\color{brown}{ 8}} x&-{\color{blue}{ 6}}=y\\ \uparrow &\uparrow \\ {\color{brown}{ slope}}&{\color{blue}{ y-intercept}} \end{array}\impliedby slope-intercept\ form\)

OpenStudy (iwanttogotostanford):

ok, thanks

OpenStudy (anonymous):

np

OpenStudy (jdoe0001):

well... actaullly includes the negative, so \(8x-y-6=0\implies \begin{array}{llll} {\color{brown}{ 8}} x&{\color{blue}{- 6}}=y\\ \uparrow &\uparrow \\ {\color{brown}{ slope}}&{\color{blue}{ y-intercept}} \end{array}\impliedby slope-intercept\ form\)